OpenStudy (anonymous):

Given the sequence 1331, 1000, 729, x, y, z, 125, . . . , what is the sum of x, y, z?

OpenStudy (phantom):

Have you recognized the fact 1331= 11^3, 1000= 10^3 and 729= 9^3.

OpenStudy (phantom):

So with 125 being 5^3 , x= 8^3=512, y= 7^3 =343 and z= 6^3 =216 if I am right with calculations :P

OpenStudy (anonymous):

Thank you so much